# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ 2. Gromov-Wasserstein (GW) Distances
[…]
@@
[…]
@@
For a simple visualization of weak isomorphic networks, see Figure 4 (and also Figure 10 described by Example B.7 in App. B). It is clear that if two networks are strongly isomorphic, then they are also weakly isomorphic. Moreover, if two networks are weakly isomorphic and have the structure of mm-spaces, then they are strongly isomorphic; see Proposition B.6 in App. B.3 for the finite case.
Figure 4. Illustration of weakly isomorphic networks $\mathbb{X}$ and $\mathbb{Y}$. The network $\mathbb{X}$ consists of one point (i.e., $X$ is a singleton) with unit mass (i.e., $\mu_{X}$ is the delta-measure), and $\omega_{X}$ represented by the $1\times 1$ matrix $\mathbf{X}=(3)$. The network $\mathbb{Y}$ is given by the probability space $(Y,\mu_{Y})$ consisting of two points with equal mass, and $\omega_{Y}$ represented by the $2\times 2$ matrix $\mathbf{Y}=\begin{pmatrix}3&3\\ 3&3\end{pmatrix}$. The pivoting space $(Z,\mu_{Z})$ in Definition 2.3 can be taken as $(Y,\mu_{Y})$, with $\varphi_{X}=\varphi$ (red assignment) and $\varphi_{Y}$ the identity map.
Consider the equivalence relation on the space of of networks given by
$\mathbb{X}\sim\mathbb{X}^{\prime}\quad\text{ if and only if }\quad GW(\mathbb{X},\mathbb{X}^{\prime})=0.$ (7)
Then, $GW(\cdot,\cdot)$ defines a metric on the quotient space of mm-spaces under this equivalence relation. Moreover, (7) coincides with the notion of strong isomorphism [38, Thm. 5.1]. The extension of the GW discrepancy from mm-spaces to general networks is not necessarily a metric modulo strong isomorphisms [13, Rem. 2.5 and Fig. 2]. However, in analogy to Gromov–Hausdorff distance [12], in [13, Thms. 2.3 and 2.4] the authors prove that GW in (6) is a pseudo-metric in the space of networks (i.e, non-negative, symmetric and satisfies the triangle inequality). Moreover, $GW(\mathbb{X},\mathbb{Y})=0$ if and only if $\mathbb{X}$ and $\mathbb{Y}$ are weakly isomorphic. As a summary, we provide the following diagram:
@@
In this paper, we consider general networks as described in Definition 2.1, up to the equivalence relation induced by weak isomorphisms. Specifically, let $\mathcal{GW}$ denote the space of equivalence classes of networks under the weak isomorphism relation $\sim^{w}$. Given a network $\mathbb{Y}=(Y,\omega_{Y},\mu_{Y})$, we denote its equivalence class by $[\mathbb{Y}]$.
@@ 2.1. GW on Finite Spaces
We follow the approach in [38, Sec. 7], where the author introduces a computationally motivated formulation of the GW problem (see also [46]).
@@
  • $\mathrm{p}_{\mathbf{X}}=(\mathrm{p}_{\mathbf{X}}[1],\dots,\mathrm{p}_{\mathbf{X}}[N_{\mathbf{X}}])\in\mathbb{R}^{N_{\mathbf{X}}}$, with $\mathrm{p}_{\mathbf{X}}[j]>0$ $\forall j\in[N_{\mathbf{X}}]$ and $\sum_{j\in[N_{\mathbf{X}}]}\mathrm{p}_{\mathbf{X}}[j]=1$, encodes the discrete measure $\mu_{X}$ supported on the $N_{\mathbf{X}}$ points; that is,
    $\mu_{X}=\sum_{j=1}^{N_{\mathbf{X}}}\mathrm{p}_{\mathbf{X}}[j]\delta_{x_{j}}.$
  • $\mathbf{X}\in\mathbb{R}^{N_{\mathbf{X}}\times N_{\mathbf{X}}}$ is a real $N_{\mathbf{X}}\times N_{\mathbf{X}}$ matrix, sometimes referred to as a dissimilarity matrix, where each entry $(i,j)$ represents the interaction between nodes $x_{i}$ and $x_{j}$; that is,
    $\omega_{X}(x_{i},x_{j})=\mathbf{X}_{ij}.$
@@
$GW((\mathbf{X},\mathrm{p}),(\mathbf{Y},\mathrm{q}))^{2}=\min_{\pi\in\Pi(\mathrm{p},\mathrm{q})}\sum_{i\in[N],j\in[M]}\sum_{k\in[N],l\in[M]}|\mathbf{X}_{ik}-\mathbf{Y}_{jl}|^{2}\pi_{ij}\pi_{kl}.$ (8)
It is a quadratic optimization problem with linear constraints, with no guarantees to be convex in general, and a minimizer is always attained [38]. When considering symmetric positive (semi)definite matrices, (8) is convex. For completeness, we provide a proof of this result in Proposition B.1 in App. B.1. However, we note that distance matrices $\mathbf{X}_{ij}=d_{X}(x_{i},x_{j})$ are not generally positive semidefinite.
@@ 3. Fixed-Point Iteration for Gromov-Wasserstein Barycenters
In this section, we begin by revisiting one of the most widely used approaches for synthesizing GW barycenters [46], but without incorporating any regularization term (3.1). We then show that this iterative scheme constitutes a fixed-point iteration (Theorem 3.3 and its Corollary 3.4). Building on this, in Section 3.2 we present our new contributions on solving the GW analysis problem, including Theorem 3.6 and Algorithm 2.
@@ 3.1. GW Barycenters via a Fixed-Point Approach: The Synthesis Problem
$\Delta_{S-1}\ni\lambda\longmapsto\mathbf{Y}_{\lambda}\in\operatorname*{arg\,min}_{\mathbf{Y}\in\mathbb{R}^{M\times M}}\sum_{s\in[S]}\lambda_{s}\,GW((\mathbf{X}^{s},\mathrm{p}^{s}),(\mathbf{Y},\mathrm{q}))^{2}.$ (9)
See Proposition B.4 in App. B.2 for a discussion on the existence of a minimum in (9). If the general synthesis problem (1) is replaced by (9), we define the following barycentric space.
@@
$=\min_{\mathbf{Y}\in\mathbb{R}^{M\times M}}\min_{\{\pi^{s}\in\Pi(\mathrm{p}^{s},\mathrm{q})\}_{s=1}^{S}}\sum_{s\in[S]}\lambda_{s}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}|\mathbf{X}_{ik}^{s}-\mathbf{Y}_{jl}|^{2}\pi_{ij}^{s}\pi_{kl}^{s}.$ (11)
As mentioned in [46], when the variables are treated separatly, this optimization problem is convex with respect to $\mathbf{Y}$, but not with respect to $\{\pi^{s}\}_{s=1}^{S}$ (it is quadratic, but the associated form is not necessarily positive semidefinite) [^5] However, if the set of couplings $\{\pi^{s}\}_{s\in[S]}$ is not fixed in advance, then each $\pi^{s}$ may depend on $\mathbf{Y}$ (since it is optimal for $GW((\mathbf{X}^{s},\mathrm{p}^{s}),(\mathbf{Y},\mathrm{q}))$), implying that (3.1) is not necessarily convex in $\mathbf{Y}$. This is due to the general fact that while the supremum of convex functions is convex, the minimum is not; that is, even if $x\mapsto h(x,y)$ is convex for each $y$, the function $g(x)=\min_{y}h(x,y)$ need not be convex.. In what follows we recall [46, Prop. 3] [^6] Although [46, Prop. 3] allows for more general loss functions in the GW problem, we restrict our attention to the standard case $|\cdot|^{2}$ for clarity and simplicity..
Proposition 3.2 (Prop. 3 [46] - Revisited proof in App. B.2).
@@
$\mathbb{R}^{M\times M}\ni\mathbf{Y}\longmapsto J_{\lambda}(\mathbf{Y}):=$ $\sum_{s\in[S]}\lambda_{s}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}|\mathbf{X}_{ik}^{s}-\mathbf{Y}_{jl}|^{2}\pi_{ij}^{s}\pi_{kl}^{s}.$ (12)
@@
$\mathbf{Y}^{*}_{jl}:=\frac{1}{\mathrm{q}_{j}\mathrm{q}_{l}}\sum_{s\in[S]}\lambda_{s}\left((\pi^{s})^{T}\mathbf{X}^{s}\pi^{s}\right)_{jl},\qquad\forall l,j\in[M].$ (13)
The following result provides a closed-form expression for the elements of the GW barycenter space in Definition 3.1.
@@
$\mathbf{Y}\longmapsto J_{\lambda}(\mathbf{Y}):=\sum_{s\in[S]}\mathbf{\lambda}_{s}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}|\mathbf{X}_{ik}^{s}-{\mathbf{Y}}_{jl}|^{2}\pi_{ij}^{*s}\pi_{kl}^{*s}.$ (14)
@@
$\rho_{\lambda}(\mathbf{Y}):=\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot\sum_{s\in[S]}\lambda_{s}\,(\pi(\mathbf{Y},{s}))^{T}\,\mathbf{X}^{s}\,\pi(\mathbf{Y},{s})\qquad\forall\,\mathbf{Y}\in\mathbb{R}^{M\times M}.$ (15)
@@
Remark 3.5 (See also Example B.5 in App. B.2).
@@ 3.2. GW Barycenters via a Fixed-Point Approach: The Analysis Problem
$\mathbb{R}^{M\times M}\ni\mathbf{Y}\longmapsto\lambda_{\mathbf{Y}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}(\mathbf{Y},\mathbf{Y}_{\lambda})^{2},$ (16)
@@
$\mathbf{Y}=\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot\sum_{s\in[S]}(\lambda_{\mathbf{Y}})_{s}(\pi^{s})^{T}\mathbf{X}^{s}\pi^{s},\qquad\text{ for some }\lambda_{\mathbf{Y}}\in\Delta_{S-1},$ (17)
where, for each $s\in[S]$, $\pi^{s}$ is an optimal plan for $GW((\mathbf{X}^{s},\mathrm{p}^{s}),(\mathbf{Y},\mathrm{q}))$. It is important to notice that each $\pi^{s}$ depends on the input $\mathbf{Y}$, but, by Remark 3.5, we could consider another set of optimal plans $\{\widetilde{\pi^{s}}\}_{s=1}^{S}$ and the expression (17) would not differ. Let us simplify our notation by defining, for each $s\in[S]$, the matrices
$F(\mathbf{Y},s):=\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot(\pi^{s})^{T}\mathbf{X}^{s}\pi^{s}\in\mathbb{R}^{M\times M}.$ (18)
This implies that problem (16) for the given $\mathbf{Y}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s=1}^{S})$ coincides with
$\lambda_{\mathbf{Y}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}^{2}\left(\mathbf{Y},\,\sum_{s\in[S]}\lambda_{s}F(\mathbf{Y},s)\right).$ (19)
In other words, we aim to minimize the functional $\Theta:\mathbb{R}^{S}\to[0,\infty)$ defined by
$\Theta(\lambda):=d_{a}^{2}\left(\mathbf{Y},\sum_{s\in[S]}\lambda_{s}F(\mathbf{Y},s)\right)$ (20)
over the convex set $\Delta_{S-1}$. As the choice of divergence $d_{a}(\cdot,\cdot)$ does not affect result (as pointed out in Remark A.2), in Theorem 3.6 we will adopt $d_{a}(\mathbf{X},\mathbf{Y})=\|\mathbf{Y}-\mathbf{X}\|_{\mathrm{Frob}}$999(which, in particular, is invariant under orthogonal transformations).
@@
Let $\{(\mathbf{X}^{s},\mathrm{p}^{s})\in\mathbb{R}^{N^{s}\times N^{s}}\times\mathcal{P}_{N^{s}}\}_{s\in[S]}$, $M\in\mathbb{N}$, $(\mathbf{Y},\mathrm{q})\in\mathbb{R}^{M\times M}\times\mathcal{P}_{M}$. Let $F(\mathbf{Y},s)$ be as in (18), for each $s\in[S]$, and consider the loss $d_{a}(\mathbf{X},\mathbf{Y})=\|\mathbf{Y}-\mathbf{X}\|_{\mathrm{Frob}}$ in the formulation of the GW analysis problem (16). If $\mathbf{Y}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]})$, then the corresponding weight vector $\lambda_{\mathbf{Y}}\in\Delta_{S-1}$ associated with $\mathbf{Y}$ is given by a solution of the following convex quadratic minimization problem:
$\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{Q}\lambda,\quad\text{where }\,\mathcal{Q}_{sr}:=\mathrm{tr}\left(\left(F(\mathbf{Y},s)-\mathbf{Y}\right)^{T}\left(F(\mathbf{Y},r)-\mathbf{Y}\right)\right)\,\,\,\forall s,r\in[S].$ (21)
@@
${K\lambda=b}\qquad\text{ subject to }\lambda\in\Delta_{S-1}\subset\mathbb{R}^{S},$ (22)
@@
$K_{sr}:=\mathrm{tr}\left(F(\mathbf{Y},s)^{T}F(\mathbf{Y},r)\right)\quad\forall s,r\in[S],\quad b_{s}:=\mathrm{tr}\left(\mathbf{Y}^{T}\,F(\mathbf{Y},s)\right)\quad\forall s\in[S].$ (23)
@@
Due to (17), the GW barycentric coordinate vector $\lambda_{\mathbf{Y}}$ is such that $\Theta(\lambda_{\mathbf{Y}})=\lambda_{\mathbf{Y}}^{T}\mathcal{Q}\lambda_{\mathbf{Y}}=0$. Equivalently, $\lambda_{\mathbf{Y}}$ is a solution of the quadratic minimization problem (21), that is,
$\lambda_{\mathbf{Y}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{Q}\lambda.$
The matrix $\mathcal{Q}$ is symmetric and positive semidefinite. This holds since it is the Gram matrix of the system of matrices $\{F(\mathbf{Y},s)-\mathbf{Y}\}_{s=1}^{S}$ with respect to the inner product in $\mathbb{R}^{M\times M}$ given by $\langle A,B\rangle=\mathrm{tr}(A^{T}B)$, and it is well-known that all Gram matrices are symmetric positive semidefinite [53]. Thus, $\Theta$ is convex and non-negative as a function over the whole $\mathbb{R}^{S}$. So, critical points of $\Theta$ are global minima. Moreover, we can decompose
@@
$\min_{\lambda\in\Delta_{S-1}}\lambda^{T}K\lambda-2\lambda^{T}b,$ (24)

### v2

@@ 2. Gromov-Wasserstein (GW) Distances
The natural hierarchy among the above structures can be expressed by the following chain of inclusions: $\text{mm-spaces}\subset\text{gm-spaces}\subset\text{networks}$.
@@
(4) $GW(\mathbb{X},\mathbb{Y})=\inf_{\pi\in\Pi(\mu_{X},\mu_{Y})}{\left(\int_{X\times Y}\int_{X\times Y}|\omega_{X}(x,x^{\prime})-\omega_{Y}(y,y^{\prime})|^{p}d\pi(x,y)d\pi(x^{\prime},y^{\prime})\right)^{\frac{1}{p}}}.$
@@
For a simple visualization of weak isomorphic networks, see Figure 3 (and also Figure 12 described by Example B.7 in Suppl. Mat. B). It is clear that if two networks are strongly isomorphic, then they are also weakly isomorphic. Moreover, if two networks are weakly isomorphic and have the structure of mm-spaces, then they are strongly isomorphic [42, 13] (see Proposition B.6 in Suppl. Mat. B.3 for the finite case).
Figure 3. Illustration of weakly isomorphic networks $\mathbb{X}$ and $\mathbb{Y}$. The network $\mathbb{X}$ consists of one point (i.e., $X$ is a singleton) with unit mass (i.e., $\mu_{X}$ is the delta-measure), and $\omega_{X}$ represented by the $1\times 1$ matrix $\mathbf{X}=(3)$. The network $\mathbb{Y}$ is given by the probability space $(Y,\mu_{Y})$ consisting of two points with equal mass, and $\omega_{Y}$ represented by the $2\times 2$ matrix $\mathbf{Y}=\left(\begin{smallmatrix}3&3\\ 3&3\end{smallmatrix}\right)$. The pivoting space $(Z,\mu_{Z})$ in Definition 2.3 can be taken as $(Y,\mu_{Y})$, with $\varphi_{X}=\varphi$ (red assignment) and $\varphi_{Y}$ the identity map.
Consider the equivalence relation on the space of networks given by $\mathbb{X}\sim\mathbb{X}^{\prime}$ if and only if $GW(\mathbb{X},\mathbb{X}^{\prime})=0.$ Then, $GW(\cdot,\cdot)$ defines a metric on the quotient space of mm-spaces under this equivalence relation. Moreover, this equivalence relation coincides with the notion of strong isomorphism [42, Thm. 5.1]. The extension of the GW discrepancy from mm-spaces to general networks is not necessarily a metric modulo strong isomorphisms [14, Rem. 2.5 and Fig. 2]. However, in analogy to Gromov-Hausdorff distance [13], in [14, Thms. 2.3 and 2.4] the authors prove that GW in (4) is a pseudo-metric in the space of networks (i.e, non-negative, symmetric and satisfies the triangle inequality). Moreover, $GW(\mathbb{X},\mathbb{Y})=0$ if and only if $\mathbb{X}$ and $\mathbb{Y}$ are weakly isomorphic. As a summary, we provide the following diagram:
@@
In this paper, we consider general networks as described in Definition 2.1, up to the equivalence relation induced by weak isomorphisms. Specifically, let $\mathcal{GW}$ denote the space of equivalence classes of networks under the weak isomorphism relation $\sim^{w}$. Given a network $\mathbb{Y}=(Y,\omega_{Y},\mu_{Y})$, we denote its equivalence class by $\mathbb{Y}]$.
@@ 2.1. GW on Finite Spaces
We follow the approach in [42, Sec. 7], where the author introduces a computationally motivated formulation of the GW problem (see also [49]).
@@
  • $\mathrm{p}_{\mathbf{X}}=(\mathrm{p}_{\mathbf{X}}[1],\dots,\mathrm{p}_{\mathbf{X}}[N_{\mathbf{X}}])\in\mathbb{R}^{N_{\mathbf{X}}}$, with $\mathrm{p}_{\mathbf{X}}[j]>0$ $\forall j\in[N_{\mathbf{X}}]$ and $\sum_{j\in[N_{\mathbf{X}}]}\mathrm{p}_{\mathbf{X}}[j]=1$, encodes the discrete measure $\mu_{X}$ supported on the $N_{\mathbf{X}}$ points; that is, $\mu_{X}=\sum_{j=1}^{N_{\mathbf{X}}}\mathrm{p}_{\mathbf{X}}[j]\delta_{x_{j}}.$
  • $\mathbf{X}\in\mathbb{R}^{N_{\mathbf{X}}\times N_{\mathbf{X}}}$ is a real $N_{\mathbf{X}}\times N_{\mathbf{X}}$ matrix, sometimes referred to as a dissimilarity matrix, where each entry $(i,j)$ represents the interaction between nodes $x_{i}$ and $x_{j}$; that is, $\omega_{X}(x_{i},x_{j})=\mathbf{X}_{ij}.$
@@
(5) $GW((\mathbf{X},\mathrm{p}),(\mathbf{Y},\mathrm{q}))^{2}=\min_{\pi\in\Pi(\mathrm{p},\mathrm{q})}\sum_{i\in[N],j\in[M]}\sum_{k\in[N],l\in[M]}|\mathbf{X}_{ik}-\mathbf{Y}_{jl}|^{2}\pi_{ij}\pi_{kl}.$
It is a quadratic optimization problem with linear constraints, with no guarantees to be convex in general, and a minimizer is always attained [42]. When considering symmetric positive (semi)definite matrices, (5) is convex. For completeness, we provide a proof of this result in Proposition B.1 in Suppl. Mat. B.1. However, we note that distance matrices $\mathbf{X}_{ij}=d_{X}(x_{i},x_{j})$ are not generally positive semidefinite.
@@ 3. Fixed-Point Iteration for Gromov-Wasserstein Barycenters
In this section, we begin by revisiting one of the most widely used approaches for synthesizing GW barycenters [49], but without incorporating any regularization term (Section 3.1). We then show that this iterative scheme constitutes a fixed-point iteration (Theorem 3.3 and its Corollary 3.4). Building on this, in Section 3.2 we present our new contributions on solving the GW analysis problem, including Theorem 3.6 and Algorithm 2.
@@ 3.1. GW Barycenters via a Fixed-Point Approach: The Synthesis Problem
(6) $\Delta_{S-1}\ni\lambda\longmapsto\mathbf{Y}_{\lambda}\in\operatorname*{arg\,min}_{\mathbf{Y}\in\mathbb{R}^{M\times M}}\sum_{s\in[S]}\lambda_{s}\,GW((\mathbf{X}^{s},\mathrm{p}^{s}),(\mathbf{Y},\mathrm{q}))^{2}.$
See Prop. B.4 in Suppl. Mat. B.2 for a discussion on the existence of a minimum in (6). If the general synthesis problem (1) is replaced by (6), we define the following barycentric space.
@@
(8) $=\min_{\mathbf{Y}\in\mathbb{R}^{M\times M}}\min_{\{\pi^{s}\in\Pi(\mathrm{p}^{s},\mathrm{q})\}_{s=1}^{S}}\sum_{s\in[S]}\lambda_{s}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}|\mathbf{X}_{ik}^{s}-\mathbf{Y}_{jl}|^{2}\pi_{ij}^{s}\pi_{kl}^{s}.$
As mentioned in [49], when the variables are treated separately, this optimization problem is convex with respect to $\mathbf{Y}$, but not with respect to $\{\pi^{s}\}_{s=1}^{S}$ (it is quadratic, but the associated form is not necessarily positive semidefinite) [^5] However, if the set of couplings $\{\pi^{s}\}_{s\in[S]}$ is not fixed in advance, then each $\pi^{s}$ may depend on $\mathbf{Y}$ (since it is optimal for $GW((\mathbf{X}^{s},\mathrm{p}^{s}),(\mathbf{Y},\mathrm{q}))$), implying that (3.1) is not necessarily convex in $\mathbf{Y}$. This is due to the general fact that while the supremum of convex functions is convex, the minimum is not; that is, even if $x\mapsto h(x,y)$ is convex for each $y$, the function $g(x)=\min_{y}h(x,y)$ need not be convex.. In what follows we recall [49, Prop. 3] [^6] Although [49, Prop. 3] allows for more general loss functions in the GW problem, we restrict our attention to the standard case $|\cdot|^{2}$ for clarity and simplicity..
Proposition 3.2 (Prop. 3 [49] - Revisited proof in Suppl. Mat. B.2).
@@
(9) $\mathbb{R}^{M\times M}\ni\mathbf{Y}\longmapsto J_{\lambda}(\mathbf{Y}):=\sum_{s\in[S]}\lambda_{s}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}|\mathbf{X}_{ik}^{s}-\mathbf{Y}_{jl}|^{2}\pi_{ij}^{s}\pi_{kl}^{s}.$
@@
(10) $\mathbf{Y}^{*}_{jl}:=\frac{1}{\mathrm{q}_{j}\mathrm{q}_{l}}\sum_{s\in[S]}\lambda_{s}\left((\pi^{s})^{T}\mathbf{X}^{s}\pi^{s}\right)_{jl},\qquad\forall l,j\in[M].$
The following result provides a closed-form expression for the elements in (7).
@@
(11) $\mathbf{Y}\longmapsto J_{\lambda}(\mathbf{Y}):=\sum_{s\in[S]}\mathbf{\lambda}_{s}\sum_{i\in[N^{s}],\,j\in[M]}\sum_{k\in[N^{s}],\,l\in[M]}|\mathbf{X}_{ik}^{s}-{\mathbf{Y}}_{jl}|^{2}\pi_{ij}^{*s}\pi_{kl}^{*s}.$
@@
(12) $\rho_{\lambda}(\mathbf{Y}):=\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot\sum_{s\in[S]}\lambda_{s}\,(\pi(\mathbf{Y},{s}))^{T}\,\mathbf{X}^{s}\,\pi(\mathbf{Y},{s})\qquad\forall\,\mathbf{Y}\in\mathbb{R}^{M\times M}.$
@@
Remark 3.5 (See also Example B.5 in Suppl. Mat. B.2).
@@ 3.2. GW Barycenters via a Fixed-Point Approach: The Analysis Problem
(13) $\mathbb{R}^{M\times M}\ni\mathbf{Y}\longmapsto\lambda_{\mathbf{Y}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}(\mathbf{Y},\mathbf{Y}_{\lambda})^{2},$
@@
(14) $\mathbf{Y}=\sum_{s\in[S]}(\lambda_{\mathbf{Y}})_{s}\,F(\mathbf{Y},s),\qquad\text{ for some }\lambda_{\mathbf{Y}}\in\Delta_{S-1},$
where, for each $s\in[S]$, one considers $\pi^{s}$ is an optimal plan for $GW((\mathbf{X}^{s},\mathrm{p}^{s}),(\mathbf{Y},\mathrm{q}))$ to define
(15) $F(\mathbf{Y},s):=\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot(\pi^{s})^{T}\mathbf{X}^{s}\pi^{s}\in\mathbb{R}^{M\times M}.$
It is important to notice that each $\pi^{s}$ depends on the input $\mathbf{Y}$, but, by Remark 3.5, we could consider another set of optimal plans $\{\widetilde{\pi^{s}}\}_{s=1}^{S}$ and the expression (14) would not differ. This implies that problem (13) for the given $\mathbf{Y}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s=1}^{S})$ coincides with
(16) $\lambda_{\mathbf{Y}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}^{2}\left(\mathbf{Y},\,\sum_{s\in[S]}\lambda_{s}F(\mathbf{Y},s)\right),$
where, by the non-negativity of $d_{a}(\cdot,\cdot)$ and by the GW-BCM hypothesis, we know in advance that for such $\mathbf{Y}$, $\min_{\lambda\in\Delta_{S-1}}d_{a}\left(\mathbf{Y},\,\sum_{s\in[S]}\lambda_{s}F(\mathbf{Y},s)\right)=0.$ In other words, we aim to minimize, over the convex set $\Delta_{S-1}$, the functional
(17) $\Theta:\mathbb{R}^{S}\to[0,\infty)\quad\text{ defined by }\qquad\Theta(\lambda):=d_{a}^{2}\left(\mathbf{Y},\sum_{s\in[S]}\lambda_{s}F(\mathbf{Y},s)\right).$
Notice that one can verify that $\mathbf{Y}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]})$ if and only if $d_{a}(\mathbf{Y},\mathbf{Y}_{\mathbf{\lambda}})=0$ for some $\mathbf{Y}_{\lambda}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]})$, for any choice of divergence $d_{a}(\cdot,\cdot)$ (see Remark A.2 for its formulation within the general BCM framework). In other words, to determine whether an element ‘strictly’ belongs to the set in (7), it suffices to use any nonnegative dissimilarity measure $d_{a}$ satisfying the identity of indiscernibles.
In this section, we have adopted the simplification that $\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]})$ is a subset of $\mathbb{R}^{M\times M}$ (since dissimilarity matrices have fixed size $M$, and the weight vector $\mathrm{q}$ is also fixed). As a consequence, because the choice of the divergence $d_{a}(\cdot,\cdot)$ on $\mathbb{R}^{M\times M}$ does not affect the final result, for the following Theorem 3.6, we adopt $d_{a}(\mathbf{X},\mathbf{Y})=\|\mathbf{Y}-\mathbf{X}\|_{\mathrm{Frob}}$. Considering this particular norm allows us to achieve our goal of providing closed-form expressions for the weights $\lambda$ (see (19) below)) under the GW-BCM hypothesis (i.e., in the exact case, when the input is assumed to be a true barycenter of the fixed templates). See also Corollary 3.7 reflecting the independence of solution $\lambda$ in Theorem 3.6 from the choice of $d_{a}$.
This follows the ideas of [8] in the context of the Wasserstein structure rather than GW. In that article, the loss function is chosen so that its gradient is easy to compute: the authors test different divergences ($L^{1}$, $L^{2}$, and $\mathrm{KL}$) instead of considering the Wasserstein distance (between histograms) as the loss function for the corresponding analysis problem. These choices still lead to optimization problems requiring gradient-based methods for the barycentric weights, and no simplified closed-form expression for $\lambda$ (as in (19) below) is provided.
@@
Let $\{(\mathbf{X}^{s},\mathrm{p}^{s})\in\mathbb{R}^{N^{s}\times N^{s}}\times\mathcal{P}_{N^{s}}\}_{s\in[S]}$, $M\in\mathbb{N}$, $(\mathbf{Y},\mathrm{q})\in\mathbb{R}^{M\times M}\times\mathcal{P}_{M}$. Let $F(\mathbf{Y},s)$ be as in (15), for each $s\in[S]$, and consider the loss $d_{a}(\mathbf{X},\mathbf{Y})=\|\mathbf{Y}-\mathbf{X}\|_{\mathrm{Frob}}$ in the formulation of the GW analysis problem (13). If $\mathbf{Y}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]})$, then a corresponding weight vector $\lambda_{\mathbf{Y}}\in\Delta_{S-1}$ associated with $\mathbf{Y}$ is given by a solution of the following convex quadratic minimization problem:
(18) $\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{Q}\lambda,\quad\text{where }\,\mathcal{Q}_{sr}:=\mathrm{tr}\left(\left(F(\mathbf{Y},s)-\mathbf{Y}\right)^{T}\left(F(\mathbf{Y},r)-\mathbf{Y}\right)\right)\,\,\,\forall s,r\in[S].$
@@
(19) ${K\lambda=b}\qquad\text{ subject to }\lambda\in\Delta_{S-1}\subset\mathbb{R}^{S},$
@@
(20) $K_{sr}:=\mathrm{tr}\left(F(\mathbf{Y},s)^{T}F(\mathbf{Y},r)\right)\quad\forall s,r\in[S],\quad b_{s}:=\mathrm{tr}\left(\mathbf{Y}^{T}\,F(\mathbf{Y},s)\right)\quad\forall s\in[S].$
@@
Due to (14), the GW barycentric coordinate vector $\lambda_{\mathbf{Y}}$ is such that
(21) $\Theta(\lambda_{\mathbf{Y}})=\lambda_{\mathbf{Y}}^{T}\mathcal{Q}\lambda_{\mathbf{Y}}=0.$
Equivalently, $\lambda_{\mathbf{Y}}$ is a solution of the quadratic minimization problem (18). The matrix $\mathcal{Q}$ is symmetric and positive semidefinite. This holds since it is the Gram matrix of the system of matrices $\{F(\mathbf{Y},s)-\mathbf{Y}\}_{s=1}^{S}$ with respect to the inner product in $\mathbb{R}^{M\times M}$ given by $\langle A,B\rangle=\mathrm{tr}(A^{T}B)$, and it is well-known that all Gram matrices are symmetric positive semidefinite [56]. Thus, $\Theta$ is convex and non-negative as a function over the whole $\mathbb{R}^{S}$. So, critical points of $\Theta$ are global minima. Moreover, we can decompose
@@
(22) $\min_{\lambda\in\Delta_{S-1}}\lambda^{T}K\lambda-2\lambda^{T}b,$